# Assessment of DKIST/VTF Capabilities for the Detection of Local Acoustic Source Wavefronts

### v1

@@ Abstract
Recent studies have demonstrated that temporal filtering can successfully identify local-acoustic-source wavefronts in radiative magnetohydrodynamic simulations of the solar photosphere. Extending this capability to observations promises new insight into the stochastic excitation of solar p-modes, the source depth distribution below the photosphere, and the dominant physical processes underlying acoustic wave excitation. Such measurements would also enable improved characterization of the complex wavefield in the lower chromosphere and open the possibility of ultra-local helioseismic diagnostics. In this work, we assess an observational strategy for the detection of local acoustic wavefronts on the Sun using the National Science Foundation’s Daniel K. Inouye Solar Telescope’s Visible Tunable Filter (DKIST/VTF). Because wavefront identification requires high spatial and temporal resolution and is limited by the small amplitudes of the wave perturbations, we focus on identifying specific wavelength combinations within spectral lines that maximize the sensitivity to the wave signal at the atmospheric heights where that signal is highest while minimizing contamination by atmospheric variability at other heights. Under the cadence and spectral resolution constraints of DKIST/VTF observations and for the particular simulated wavefront we examine, this approach suggests two possible strategies for the detection of acoustic wavefronts in solar observations: fast monochromatic imaging at 6302.425 Å, or ordered interleaved observations in the blue wing of either the Fe I 6302.5 Å or Fe I 5250.6 Å line (between 6302.419 Å and 6302.465 Å, or between 5250.579 Å and 5250.607 Å respectively).
@@ 3 Introduction
There are two broad classes of stellar oscillations: high amplitude pulsations and low amplitude oscillations. Self-excitation by overstability is thought to account for the former in stars such as classical Cephieds, RR Lyrae stars, $\delta$ Scuti variables, and pulsating white dwarfs, among others (Cox, 2002; Aerts et al., 2010; Daszyńska-Daszkiewicz, 2015; Samadi et al., 2015). By contrast, low-amplitude oscillations are typically associated with pulsationally stable stars, such as the Sun. Solar acoustic oscillations ($p$-modes) have radial velocity amplitudes in the photosphere of less than 10 – 20 $\rm{cm\ s^{-1}}$ with associated intensity fluctuations of a few parts per million (e.g., Christensen-Dalsgaard, 2002). No known mechanism can limit overstable pulsational modes to such low amplitudes, and so continuous stochastic excitation is thought to sustain these oscillations (Goldreich and Keeley, 1977).
@@
Differing theoretical treatments emphasize the importance of different aspects of the underlying compressible turbulence to stochastic excitation. Convective motions may transfer energy into the acoustic wave spectrum via Reynolds stresses (e.g., Lighthill, 1952, 1954; Stein, 1967; Goldreich and Keeley, 1977; Goldreich and Kumar, 1990; Balmforth, 1992; Kitiashvili et al., 2019), entropy fluctuations, (e.g., Stein and Nordlund, 1991; Goldreich et al., 1994; Rast, 1997; Nordlund and Stein, 1998; Rast, 1999; Samadi and Goupil, 2001; Samadi et al., 2001), or some combination of the two. More than one mechanism may be important, with the phase relationships between them reflecting the evolution of solar granulation (Rast, 1999). In particular, both modeling (e.g., Skartlien et al., 2000; Lindsey and Rempel, 2020) and observational (Rimmele et al., 1995; Goode et al., 1998; Strous et al., 2000; Roth et al., 2010; Bello González et al., 2010; Lindsey and Donea, 2013) evidence suggests that acoustic excitation on the Sun occurs preferentially in granular downflow lanes and is associated with strong impulsive downflow events. These may occur as a result of new downflow formation during granule fragmentation (Rast et al., 1993; Rast, 1995) or by the rapid enhancement of existing downflows. Acoustic emission in both cases includes dynamically-induced entropy fluctuations and Reynolds stresses  (Stein and Nordlund, 1991; Goldreich et al., 1994; Rast, 1997; Nordlund and Stein, 1998; Rast, 1999; Samadi and Goupil, 2001; Samadi et al., 2001). In the case of granule fragmentation, the role of entropy fluctuations is enhanced by strong localized cooling (Rast, 1997, 1999). While uncertainties remain, due to the possible presence of a correlated noise component in the spectra (Roxburgh and Vorontsov, 1997; Nigam et al., 1998, but see Philidet et al. 2020), these excitation mechanisms are broadly consistent with constraints inferred from phase-difference spectra. In particular, the measured velocity–intensity phase relationships agree with the acoustic spectral signatures expected from rapidly forming downflow plumes (Skartlien and Rast, 2000; Severino et al., 2001). Other possible mechanisms for global $p$-mode excitation include solar flares, which can trigger transient acoustic waves (Kosovichev and Zharkova, 1998; Ambastha et al., 2003; Donea and Lindsey, 2005; Martínez et al., 2020; Lindsey et al., 2020), with coupling to solar $p$-modes somewhat uncertain but plausible (Lindsey et al., 2014; Hansen et al., 2016).
@@
Thus, multiple acoustic excitation mechanisms may be active on the Sun and in other stars. Observational access to individual source events would allow for greater constraints to be placed on the source depth distributions, temporal phasing of the thermodynamic and dynamic fluctuations associated with the sources, and the coupling efficiencies to the global eigenmodes. Linking the observed wave field to specific generation sites would clarify distinct excitation pathways and their relative contributions, enable improved characterization of the complex wavefield propagating upward from the photosphere into the lower chromosphere, and open the possibility of ultra-local helioseismic diagnostics based on the local wave field emitted by the discrete sources.
@@ 3.1 The aim of this paper
In previous work, we have developed a diagnostic differencing scheme designed to isolate the small-amplitude acoustic perturbations generated by localized excitation events (Bahauddin and Rast, 2021, 2023), and have applied the method to numerical simulations to aid in the interpretation of early DKIST observations of transient wave phenomena (Bahauddin et al., 2024). The method allows identification of coherent wavefronts by leveraging the difference between the timescales of granulation evolution and that of wavefront passage across the spatial grid. It was formulated and validated using a three-dimensional numerical simulation of quiet-sun granulation (Rempel, 2014), operating directly on the temperature, pressure, and velocity fields computed by the magnetohydrodynamic simulation. The reliability of the differencing scheme thus, at face value, depends critically on having access to one or more of these physical quantities with high fidelity, and direct application to observations would require inferring the three-dimensional thermodynamic structure of the solar atmosphere from spectroscopic or spectropolarimetric measurements through inversion codes (e.g., Del Toro Iniesta and Ruiz Cobo, 1996; del Toro Iniesta and Ruiz Cobo, 2016). In practice, however, this direct approach is limited by photon noise, finite spectral, spatial, and temporal resolution, radiative transfer degeneracies, and the regularization inherent to inversion procedures, all of which preferentially suppress small-scale structure and rapidly evolving perturbations, including the wavefronts of interest here.
@@
Instead, we explore an alternative strategy that does not aim to reconstruct the full atmospheric state, but focuses on identifying the wavelengths in commonly observed photospheric lines at which the acoustic wavefront signal is most effectively captured. Using a numerical simulation, in which propagating wavefronts have already been identified, we perform spectral synthesis to determine the wavelengths at which wavefront-induced intensity fluctuations are maximum relative to background variability. Spectral response functions provide a natural quantitative framework for this analysis, as they describe how perturbations at different heights in the atmosphere map into wavelength-dependent intensity variations. Although acoustic waves are traditionally characterized through pressure perturbations in simulations and Doppler velocity measurements in observations, the detectability of these weak transient signals in spectra is primarily governed by the temperature sensitivity of a spectral line. Accordingly, our analysis focuses on temperature response functions, which directly quantify how depth-dependent thermal perturbations associated with the acoustic wavefronts translate into observable intensity variations. Doppler velocity is used only for wavefront identification and validation within the simulation, in which it is an easily accessible variable with a relatively large amplitude, while the observational strategy is optimized entirely in terms of intensity for which the response to velocity perturbations is much smaller than that to temperature fluctuations. By evaluating the response of the synthesized spectra to the height-dependent temperature fluctuations induced by a wavefront passage, we construct a detectability metric that quantifies the relative contributions of the wave signal and the background and identifies the spectral regions most favorable for wavefront detection using the successive-differencing technique.
@@
Detection and characterization of local source wavefronts places stringent and often competing requirements on spatial resolution, temporal cadence, and spectral fidelity. Acoustic wavefronts associated with localized excitation events evolve on time scales of seconds, have spatial scales of tens of kilometers, and are most visible in the simulations as they propagate across a granule. Thus, observing them with spectroscopic raster-scanning is very likely not possible. Any sequential sampling in physical space mixes temporal evolution with spatial structure, effectively smearing or suppressing the signatures of interest. Integral-field spectroscopy or fast Fabry–Pérot-based spectral line scanning (at a limited number of wavelengths) is therefore preferred, as these preserve the instantaneous spatial coherence of the wavefront while retaining spectral sensitivity. In the latter case, finite time is required to switch between wavelengths; therefore, the image time series at different wavelengths, constructed by sequentially stepping through a limited set of line positions (and thus atmospheric heights), are not strictly co-temporal. We anticipate (and discuss further in Sections 4 and 5.4) that since the wavefront perturbations in which we are interested propagate upward in the atmosphere, this temporal offset can be accounted for, and used beneficially, in observation and data analysis strategies (Section 6). Because of its combination of ultra-high spatial and spectral resolution, rapid cadence, and full-field spatial sampling, we suggest that the National Science Foundation’s Daniel K. Inouye Solar Telescope (DKIST, Rimmele et al., 2020) Visible Tunable Filter (VTF, Schmidt et al., 2016) is well positioned to meet the observational requirements of this problem (Section 6).
@@ 4 Wavefront detection
Figure 1: Propagation of a localized acoustic wavefront through the photosphere in a radiative MHD simulation. Panels show snapshots of the (A) line-of-sight velocity (top), its (B) third-order temporal difference (middle), and the corresponding (C) third-order temporal difference of temperature (bottom), revealing a coherent wavefront expanding across a granulalation cell. Temporal differencing suppresses background convective motions and isolates the expanding wavefront as it traverses a granule, demonstrating the basis of the detection technique used throughout this work. Temporal increment between image frame is 12 s. The original MURaM data cubes $(N_{x},N_{y},N_{z})=(384,384,256)$ are cropped to a $100\times 100\times 100$ pixel sub-region ($1.6\times 1.6$ Mm horizontally). The sub-region shown in this image is centered at pixel $(x_{c},y_{c})=(48,119)$, where $(x_{c},y_{c})$ denote global indices in the full MURaM cube.
Our analysis is based on a MURaM small-scale dynamo radiative magnetohydrodynamic simulation of the upper solar convection zone, photosphere, and lower chromosphere (simulation O16b SSD in Rempel, 2014). The computational domain spans $6\times 6$ Mm horizontally and 3 Mm vertically, with a grid spacing of 16 km and a cadence of 2 s. We examine indetail and accoustic wavefron that occure in the period between 1584–1682 s MURaM time.
@@
The spatial and temporal resolution of the simulation are sufficient to resolve propagating acoustic wavefronts produced by local sources. It is the same simulation that was used in the development of the wavefront detection techniques we employ (Bahauddin and Rast, 2021, 2023). The wavefront we analyze in detail in this paper is shown in Figure 1 as it propagates through the photosphere. Of particular note is the visibility of the wavefront after application of a triple temporal difference filter to the Doppler velocity ($v^{(3)}$, second row) and temperature field ($T^{(3)}$, bottom row). Wavefronts identified in the MURaM simulation typically originate from shallow sources, located tens of kilometers beneath the photosphere. As they pass through the photosphere, their horizontal phase speeds are initially very high and decrease with time as they expand, to asymptotically approach the sound speed at the source depth (Bahauddin and Rast, 2023). Once identified in velocity difference images, the propagating wavefronts are often just barely visible without temporal differencing in animated time series of the raw Doppler or temperature perturbation maps.
However, the visibility of the wavefronts is not uniform with height in the simulated atmosphere because the amplitude of the wavefront relative to the background motions (which we consider here as noise) varies with height. Figure 2 illustrates the background motions in which the wavefronts are embedded, showing the vertical velocity (A), the vertical velocity temporal difference (B), and the temperature fluctuations about the horizontal mean (C), as functions of height at a single instant of time. The region of convective overshoot, here taken as $z\sim 0.0$ to $0.42$ Mm above the photosphere (delimited with horizontal black dashed fiducial lines), has the lowest amplitude background noise. Convective amplitudes increase below the photosphere ($z\lesssim 0.0$), and the domain is filled with large amplitude acoustic gravity waves above $z\sim 0.42$. If the simulation accurately represents the Sun, observations aiming to uncover impulsively generated wavefronts have the highest likelihood of success if they are focused on the solar convective overshoot region. Since the domain is quite shallow, the upper boundary condition may play some role in establishing the chromospheric wave field. However, a wave field of similar amplitude is also found in deeper simulations (those of Rempel, 2025) with an upper boundary that extends into the corona. The origin, persistence, and importance of this wavefield to chromospheric structure is being pursued.

### v2

@@ Abstract
Recent studies have demonstrated that temporal filtering can successfully identify local-acoustic-source wavefronts in radiative magnetohydrodynamic simulations of the solar photosphere. Extending this capability to observations promises new insight into the stochastic excitation of solar p-modes, the source depth distribution below the photosphere, and the dominant physical processes underlying acoustic wave excitation. Such measurements would also enable improved characterization of the complex wavefield in the lower chromosphere and open the possibility of ultra-local helioseismic diagnostics. In this work, we assess an observational strategy for the detection of local acoustic wavefronts on the Sun using the Visible Tunable Filter (VTF) instrument on the National Science Foundation’s Daniel K. Inouye Solar Telescope (DKIST). Because wavefront identification requires high spatial and temporal resolution and is limited by the small amplitudes of the wave perturbations, we focus on identifying specific wavelength combinations within spectral lines that maximize the sensitivity to the wave signal. Under the cadence and spectral resolution constraints of DKIST/VTF observations and for the particular simulated wavefront we examine, this approach suggests two possible strategies: fast monochromatic imaging at 6302.425 Å, or ordered interleaved observations in the blue wing of either the Fe I 6302.5 Å or Fe I 5250.6 Å line (between 6302.419 Å and 6302.465 Å, or between 5250.579 Å and 5250.607 Å respectively).
@@ 3 Introduction
[…]
@@
Differing theoretical treatments emphasize the importance of different aspects of the underlying compressible turbulence to stochastic excitation. Convective motions may transfer energy into the acoustic wave spectrum via Reynolds stresses (e.g., Lighthill, 1952, 1954; Stein, 1967; Goldreich and Keeley, 1977; Goldreich and Kumar, 1990; Balmforth, 1992; Kitiashvili et al., 2019), entropy fluctuations, (e.g., Stein and Nordlund, 1991; Goldreich et al., 1994; Rast, 1997; Nordlund and Stein, 1998; Rast, 1999; Samadi and Goupil, 2001; Samadi et al., 2001), or some combination of the two. More than one mechanism may be important, with the phase relationships between them possibly reflecting the evolution of solar granulation as the acoustic source (Rast, 1999). In particular, both modeling (e.g., Skartlien et al., 2000; Lindsey and Rempel, 2020) and observational (Rimmele et al., 1995; Goode et al., 1998; Strous et al., 2000; Roth et al., 2010; Bello González et al., 2010; Lindsey and Donea, 2013) evidence suggests that acoustic excitation on the Sun occurs preferentially in granular downflow lanes and is associated with strong impulsive downflow events. These may occur as a result of new downflow formation during granule fragmentation (Rast et al., 1993; Rast, 1995) or by the rapid enhancement of existing downflows.
Acoustic emission in both cases includes dynamically-induced entropy fluctuations and Reynolds stresses  (Stein and Nordlund, 1991; Goldreich et al., 1994; Rast, 1997; Nordlund and Stein, 1998; Rast, 1999; Samadi and Goupil, 2001; Samadi et al., 2001). In the case of granule fragmentation, the role of entropy fluctuations is enhanced by strong localized cooling (Rast, 1997, 1999). Although uncertainties remain due to the possible presence of a correlated noise component in the spectra (Roxburgh and Vorontsov, 1997; Nigam et al., 1998, but see Philidet et al. 2020), these excitation mechanisms are broadly consistent with constraints inferred from helioseismic phase-difference spectra. In particular, the measured velocity–intensity phase relationships agree with the spectral signatures expected from rapidly forming downflow plumes (Skartlien and Rast, 2000; Severino et al., 2001). Other possible mechanisms for global $p$-mode excitation include solar flares, which can trigger transient acoustic waves (Kosovichev and Zharkova, 1998; Ambastha et al., 2003; Donea and Lindsey, 2005; Martínez et al., 2020; Lindsey et al., 2020), with coupling to solar $p$-modes somewhat uncertain but plausible (Lindsey et al., 2014; Hansen et al., 2016).
@@
Thus, multiple acoustic excitation mechanisms may be active on the Sun and in other stars. Observational access to individual source events (beyond those caused by flares) would allow for greater constraints to be placed on the turbulent source depth distributions, the temporal phasing between the thermodynamic and dynamic fluctuations associated with the sources, and the coupling efficiencies to the global eigenmodes. Linking the observed wave field to specific generation sites would clarify distinct excitation pathways and their relative contributions, enable improved characterization of the complex wavefield propagating upward from the photosphere into the lower chromosphere, and open the possibility of ultra-local helioseismic diagnostics based on the local wave field emitted by the discrete sources. While current local helioseismic techniques leverage primarily the local wavefield of the global p-modes (e.g., Gizon and Birch, 2005), the focus here is on exploiting the near-field of individual sources to deduce subsurface properties on granular and subgranular scales. As an example, this would allow for the diagnosis of magnetic field production during granule fragmentation (Rempel, 2018).
@@ 3.1 The aim of this paper
In previous work, we have developed a diagnostic temporal differencing method designed to isolate small-amplitude acoustic perturbations generated by localized excitation events (Bahauddin and Rast, 2021, 2023). We subsequently applied this method to numerical simulations to aid in the interpretation of early DKIST observations of transient wave phenomena (Bahauddin et al., 2024). Based originally on a neural network designed to identify the local wave response to a discrete source, the method employs a third-order forward temporal finite-difference operator that, due to the difference in time scales, both suppresses the background granulation contributions and amplifies the signal of the rapid wavefront passage across the spatial grid.
@@
The method was formulated and validated using a three-dimensional radiative magnetohydrodynamical simulation of quiet-Sun granulation (Rempel, 2014), operating directly on the temperature, pressure, and velocity fields computed by the simulation. The reliability of the differencing method, at face value, thus depends critically on having access to one or more of these physical quantities with high fidelity. Direct application to observations would require inferring the three-dimensional thermodynamic structure of the solar atmosphere from spectroscopic or spectropolarimetric measurements through inversion codes (e.g., Del Toro Iniesta and Ruiz Cobo, 1996; del Toro Iniesta and Ruiz Cobo, 2016). In practice, however, this direct approach is limited by photon noise, finite spectral, spatial, and temporal resolution, radiative transfer degeneracies, and the regularization inherent to inversion procedures, all of which preferentially suppress small-scale structure and rapidly evolving perturbations, including the wavefronts of interest here. Instead, we explore an alternative strategy that does not aim to reconstruct the full atmospheric state but focuses on identifying the wavelengths in commonly observed photospheric lines at which the acoustic wavefront signal is most effectively captured.
@@
Using the same numerical simulation in which we have already identified propagating wavefronts, we perform spectral synthesis to determine the wavelengths at which wavefront-induced intensity fluctuations are maximized relative to background variability. Spectral response functions provide a natural quantitative framework for this analysis, as they describe how perturbations at different heights in the atmosphere map into wavelength-dependent intensity variations. Although acoustic waves are traditionally characterized through pressure perturbations in simulations and Doppler velocity measurements in observations, the detectability of these weak transient signals in spectra is primarily governed by the temperature sensitivity of a spectral line. Accordingly, our analysis focuses on the temperature response functions, which directly quantify how depth-dependent thermal perturbations associated with the acoustic wavefronts translate into observable intensity variations. By evaluating the response of the synthesized spectra to the height-dependent temperature fluctuations induced by a wavefront passage, we construct a detectability metric that quantifies the relative contributions of the wave signal and of the background and identifies the spectral regions most favorable for wavefront detection using the successive temporal differencing technique.
@@
Detection and characterization of local source wavefronts places stringent and often competing requirements on spatial resolution, temporal cadence, and spectral fidelity. Acoustic wavefronts associated with localized excitation events evolve on time scales of seconds, have spatial scales of tens of kilometers, and are most visible in the simulations as they propagate across a granule. Thus, observing them with spectroscopic raster-scanning is very likely not possible. Any sequential sampling in physical space mixes temporal evolution with spatial structure, effectively smearing or suppressing the signatures of interest. Integral-field spectroscopy or fast Fabry–Pérot-based spectral line scanning (at a limited number of wavelengths) is therefore preferred, as these preserve the instantaneous spatial coherence of the wavefront while retaining spectral sensitivity. In the latter case, finite time is required to switch between wavelengths, and the image time series at different wavelengths, constructed by sequentially stepping through a limited set of line positions (and thus atmospheric heights), are not strictly co-temporal. We anticipate (and discuss further in Sections 4 and 5.4) that since the wavefront perturbations in which we are interested propagate upward in the atmosphere, this temporal offset can be accounted for, and used beneficially, in observation and data analysis strategies (Section 6). Because of its combination of ultra-high spatial and spectral resolution, rapid cadence, and full-field spatial sampling, we suggest that the National Science Foundation’s Daniel K. Inouye Solar Telescope (DKIST, Rimmele et al., 2020) Visible Tunable Filter (VTF, Schmidt et al., 2016; Kentischer et al., 2025) is well positioned to meet the observational requirements of this problem (Section 6).
@@ 4 Wavefront detection
Figure 1: Propagation of a localized acoustic wavefront through the photosphere in a radiative MHD simulation. Panels show snapshots of the (A) line-of-sight velocity (top), its (B) third-order temporal difference (middle), and the corresponding (C) third-order temporal difference of temperature (bottom), revealing a coherent wavefront expanding across a granulation cell. Temporal differencing suppresses background convective motions and isolates the expanding wavefront as it traverses a granule, demonstrating the basis of the detection technique used throughout this work. Plotting cadence (sub-sampled for display) is 12 s, while the temporal difference is applied to the native simulation cadence of 2 s. The original MURaM data cubes $(N_{x},N_{y},N_{z})=(384,384,256)$ are cropped to a $100\times 100\times 100$ pixel sub-region ($1.6\times 1.6$ Mm horizontally). The sub-region shown in this image is centered at pixel $(x_{c},y_{c})=(48,119)$, where $(x_{c},y_{c})$ denote global indices in the full MURaM cube.
Our analysis is based on a MURaM quiet-Sun, mixed-polarity small-scale dynamo radiative magnetohydrodynamic (MHD) simulation of the upper solar convection zone, photosphere, and lower chromosphere ($O16b$ in Rempel, 2014). The computational domain of the simulation spans 1 hour of solar time over a $6\times 6\times 3$ Mm field of view with a grid spacing of 16 km and a 2 s cadence. This spatial and temporal resolution is sufficient to resolve propagating acoustic wavefronts produced by local sources. In the simulated photosphere, weak, small-scale magnetic fields are concentrated in intergranular lanes with nearby opposite polarities, with no coherent magnetic structure on the scale of plage. In the upper layers of the domain, the field becomes increasingly horizontally inclined, and the targeted photospheric wavefront signatures are primarily affected by variability in the convective background not the magnetic field.
In previous work, we identified 25 wavefronts of sufficient diagnostic quality (i.e., semicircular outwardly propagating perturbations visible over at least 6 or 7 timesteps) in the simulation (Bahauddin and Rast, 2023). These typically originate from shallow sources located some tens of kilometers beneath the photosphere. As the wavefronts pass through the photosphere, their horizontal phase speeds are initially very high and decrease with time as they expand, to asymptotically approach the sound speed at the source depth (Bahauddin and Rast, 2023). Once identified in velocity difference images, the propagating wavefronts are often just barely visible without temporal differencing in animated time series of the raw Doppler or temperature perturbation maps.
@@
The wavefront we analyze in detail in this paper occurs between 1584–1682 s in the simulation and is the same event analyzed in Bahauddin and Rast (2023). Figure 1 shows the wavefront propagating through the photosphere. Its visibility is particularly evident after applying the temporal differencing method to the Doppler velocity ($v^{(3)}$, second row) and the temperature field ($T^{(3)}$, bottom row). However, the visibility of the wavefront is not uniform with height in the simulated atmosphere because the amplitude of the wavefront relative to the background convective motions (which we consider here as noise) varies with height.
Figure 2 illustrates the background motions in which the wavefront is embedded, showing the vertical velocity (A), the vertical velocity temporal difference (B), and the temperature fluctuations about the horizontal mean (C), as functions of height at a single instant of time. The region of convective overshoot, here taken as $z\sim 0.0$ to $0.42$ Mm above the photosphere (delimited with horizontal black dashed fiducial lines), has the lowest amplitude background noise. Convective amplitudes increase below the photosphere ($z\lesssim 0.0$), and the domain is filled with large amplitude acoustic gravity waves above $z\sim 0.42$.
Time series observations of high-resolution photospheric spectra in quiet-Sun regions (e.g., Puschmann et al., 2005) show similar overshoot signatures extending through the lower photosphere, with the granule-lane temperature contrast becoming small around $\sim$0.15-0.17 Mm and vertical velocity structures remaining detectable up to heights of order $\sim$0.37 Mm. Therefore, assuming the simulation in this respect accurately represents the Sun, observations aiming to uncover impulsively generated wavefronts have the highest likelihood of success if they are focused on the solar convective overshoot region. We note that, since the simulation domain is quite shallow, the upper boundary condition may play some role in establishing the chromospheric wave field. However, a wave field of similar amplitude is also found in deeper simulations (those of Rempel, 2025) which have an upper boundary that extends into the corona. The origin, persistence, and importance of this wavefield to the chromospheric structure and dynamics are being pursued but are not the focus of this paper.